# Cross-Domain Few-Shot Learning for Hyperspectral Image Classification Based on Mixup Foundation Model

### v1

@@ Abstract
Cross-Domain Few-Shot Learning, Few-Shot Learning, Meta-learning, Transfer Learning, Hyperspectral Image.
I Introduction
@@
AHyperspectral image (HSI) plays an important role in the remote sensing (RS) field because it is, unlike RGB images, in the form of 3D data cubes and contains abundant spatial and spectral information revealing unique characteristics of various land covers [20]. It is widely applied across different application domains such as mining, biology, and natural disasters, etc.
@@
The challenge of the HSI classification task is to assign a label to each pixel, which incurs high labeling costs. Early approaches to HSI classification problems make use of traditional machine learning methods [40, 32] combined with manual feature extraction techniques. Nonetheless, such approaches are over-dependent on prior knowledge and suffer from significant bottlenecks in handling high-dimensional data such as HSIs. The advent of deep learning (DL) has revolutionized the RS field and has been successful across different RS problems, including the HSI classification problems. In contrast to traditional approaches, the DL methods feature an automatic feature engineering process, making it possible for deep and nonlinear features to be automatically extracted, learned, and classified [4]. Various models have been proposed to cope with the HSI classification problems, including CNN [21, 66, 61], transformer [14, 48], graph neural network [27, 8], etc. Nonetheless, the success of DL approaches is attributed to the availability of large datasets hindering their practical applications in the real-world HSI classification problems, imposing prohibitive labeling costs [35].
@@
Such an issue has motivated the development of various methods to reduce the labeling cost problem. The semi-supervised learning (SSL) [56, 53] and self-supervised learning (SSL) [25, 29] methods have been proposed in the realm of the HSI classification problems. Although these approaches are capable of reducing the labeling cost to some extent, they call for a lot of samples. It is not well-suited to the HSI classification problem, where the cost of data collection is also expensive. This issue has led to the development of few-shot learning methods [44, 45] for the HSI classification problems. The few-shot learning approach relies on an episodic learning procedure based on the so-called meta-learning approach. That is, the training process is divided into a number of episodes having very few samples, i.e., support and query sets. Such approaches allow to quickly generalize to unseen classes with limited samples because a model is trained to acquire meta-knowledge. Nevertheless, conventional few-shot learning methods aren’t prepared to address the domain shift problem where the source and target domains are drawn from different distributions. This bottleneck is not in line with the fact that data collection of HSIs might be influenced by external factors, e.g., weather, lighting, sensor nonlinearities, etc., causing the issue of domain shifts.
@@
Cross-scene HSI classification methods [55, 30, 19] are proposed to address the domain shift problem. This approach transfers knowledge of a label-rich domain known as the source domain to a label-scarce domain known as the target domain. The underlying assumption is that there are no labels at all in the target domain, and the presence of domain shifts between the source domain and the target domain. Usually, such approaches are driven by the domain adaptation (DA) method [18] to align the distributional discrepancies between the source domain and the target domain. The key limitation of the cross-scene HSI classification lies in the closed-set assumption, where the source domain and the target domain share the same label space. Cross-Domain Few-Shot Learning (CDFSL) [65, 33] aims to overcome this limitation, where the goal is to cope with the issue of data scarcity and domain shift simultaneously. It goes one step ahead of cross-scene HSI classification method,s where the label space between the source domain and the target domain can be completely different.
@@
The CDFSL [33] for HSI classification has grown rapidly, where numerous works have been proposed to deal with this problem. Existing works can be categorized into three groups: discrepancy-based methods [22, 15, 23, 6, 37, 65, 38, 10], adversarial-based methods [36, 43, 37, 46, 62, 57, 2, 51, 26, 42, 70] and contrastive-based methods [41, 58, 17, 63, 25]. The discrepancy-based methods are designed to minimize distributional discrepancy between the source domain and the target domain using the discrepancy metrics. The adversarial-based methods utilize the adversarial domain adaptation technique [12], playing the minmax game between the discriminator and generator. The contrastive-based methods apply the idea of contrastive learning to guarantee the compactness of intra-class data and the separability of inter-class data. Nonetheless, existing methods suffer from at least three drawbacks:
@@
  • Existing methods apply unrealistic data augmentation procedures to enlarge the size of the target domain [51]. Such an approach greatly simplifies the issue of data scarcity but may be trapped in the out-of-distribution problem. The model performance may be compromised by the problem of sample diversity. In addition, the data augmentation procedure simply follows the RGB image augmentation strategy. The RGB image augmentation strategy might not be compatible with the HSI case.
@@
  • Very few works have explored the strength of hyperspectral foundation models in the realm of CDFSL for HSI classification problems. The foundation model possesses strong generalization potential, which can be quickly adapted to downstream tasks, thus aligning very well with the spirit of the CDFSL for HSI classification problems. To the best of our knowledge, only [3] has pioneered the use of a foundation model in the context of CDFSL for HSI classification. However, this method still utilizes the foundation model of the RGB images, hindering their generalization for HSIs. Besides, we propose a new parameter-efficient fine-tuning (PEFT) method, namely coalescent projection (CP), whereas they [3] are still based on the LoRA method.
This paper proposes a mixup foundation model (MIFOMO) for CDFSL of HSI classification problems. MIFOMO is built upon a hyper-spectral foundation model, HyperSIGMA [47], pretrained in the HyperGlobal-450K dataset containing 450 K HSIs. This goes one step ahead of [3], still relying on the foundation model of RGB images, dinoV2. To remedy the issue of overfitting, we put forward a novel PEFT technique, namely CP [34]. That is, only CPs are trainable during the training process, leaving the backbone network frozen. This strategy also allows preservation of the strong generalization power of the foundation model. CP can be seen as a successor of plain prompt [52], incurring significantly fewer parameters because it relies on a single learnable matrix. In addition, the CP technique is simpler to apply than the plain prompt because there is no need to select the prompt length. Another innovation of MIFOMO lies in the domain adaptation strategy, where the mixup domain adaptation approach [11] is put into perspective. This strategy consists of two phases: source domain and intermediate domain, in which the episodic-based meta-learning approach is undertaken. Unlike direct knowledge transfer between the source and target domains, the intermediate domain functions as a bridge between the source and target domains, making seamless knowledge transfer possible because direct domain alignment may be too intricate to perform. Last but not least, the concept of label smoothing is incorporated to combat the noisy pseudo-label problem. That is, pseudo-labels are assigned to the query set of the target domain to prevent the label leakage problem.
@@
    This paper proposes the mixup foundation model (MIFOMO) for the CDFSL of HSI classification problems. This approach is unique in respect to state-of-the-art (SOTA) methods because it relies on the PEFT approach, i.e., existing works are based on the discrepancy-based methods, the adversarial-based methods, and the contrastive-based methods. It distinguishes itself from [3] because of the hyperspectral foundation model and the CP technique.
@@
    The concept of CP is introduced to adapt the foundation model, namely HyperSIGMA, to the target domain. CP uses a single trainable matrix connecting the key and query projections in the attention map while leaving the backbone network frozen. CP offers very few trainable parameters and is easy to apply because it isn’t dependent on any hyperparameters. We extend [34] designed for natural images to hyperspectral images here.
@@
    MIFOMO advances the mixup domain adaptation [11] originally designed for unsupervised domain adaptation (UDA) scenarios, i.e., our concern is in the CDFSL of the HSI classification problem here. Moreover, the mixup domain adaptation method is generalized here using the idea of label smoothing to combat the noisy pseudo-label problem when meta-learning the target domain, while the episodic-based meta-learning approach is implemented in the source and intermediate domains. Besides, [11] still works on the traditional architecture, whereas we implement such an idea in the context of a foundation model.
@@ II Related Works
Cross-Domain Few-Shot Learning (CDFSL) for the Hyperspectral Image (HSI) classification problem is meant to address the problem of data scarcity and domain shift in the HSI classification problem. It aims to leverage discriminative information of the label-rich domain (source domain) to solve the HSI classification problem within the label-scarce domain (target domain). Compared to cross-scene HSI classification problems [55, 30, 19],this problem features extreme domain discrepancies where the label spaces between the two domains might be completely disjoint. The CDFSL for HSI classification problems have grown rapidly where existing approaches can be divided into three categories based on how the domain alignment approach is performed: discrepancy-based methods [22, 15, 23, 6, 37, 65, 38, 10], adversarial-based methods [36, 43, 37, 46, 62, 57, 2, 51, 26, 42, 70] and contrastive-based methods [41, 58, 17, 63, 25].
II-A Discrepancy-based method
@@
The discrepancy-based method measures the domain gap and minimizes it via a distance metric. [22] utilizes the Wasserstein distance metric, [15] applies the class covariance matrix, [23] uses the semantic information as complementary information of prototype construction, [6] implements the CORAL strategy, [37] makes use of the MMD metric. [65] applies the graph mining strategy, [38] designs the multiorder spectral interaction block, [10] uses the knowledge distillation strategy. The aforementioned works directly minimize the discrepancy between two domains, which can be difficult in the case of extreme domain gaps and often result in negative transfer. Our approach here is based on the mixup domain adaptation, consisting of two stages: the source domain and the intermediate domain. The intermediate domain is crafted through the mixup mechanism and allows smooth knowledge transfer because it functions as a bridge between the source domain and the target domain. In addition, the direct domain adaptation method performs global domain alignments, ignoring the class alignment step, i.e., local. Note that the label spaces of the two domains don’t overlap.
II-B Adversarial-based Method
@@
The adversarial-based method follows an adversarial domain adaptation strategy [12] which introduces a domain discriminator detecting the origin of data samples, i.e., source or target domains. It plays a minmax game with the feature extractor, and the domain alignment stage is achieved when the feature extractor successfully fools the domain discriminator. [36] applies the vanilla adversarial domain adaptation method, [43] implements the conditional adversarial domain adaptation method, [37] constructs an intermediate domain, [46] devises the local and global adversarial domain adaptation method, [62] utilizes the collaborative adversarial network concept. [57] inserts a varying weight to the conditional adversarial loss, [2] uses a semantic-aware strategy for the domain discriminator, [51] implements the random sampling strategy to address the curse of dimensionality, and [26] applies the prototype-based conditional adversarial domain adaptation method. [42] proposes a cyclic domain triplet loss combining the adversarial domain adaptation technique and the contrastive learning strategy, while [70] offers a cyclic resemblance domain adversarial network. As with the discrepancy-based method, the adversarial-based method performs the global domain alignment strategy, excluding the class alignment step.
II-C Contrastive-based method
@@
The contrastive-based method relies on the contrastive learning strategy to attain alignments between the two domains. [41] proposes the inter-domain contrastive loss function, [58] puts forward the graph convolutional contrast module, [17] puts into perspective the kernel triplet loss function, [63, 25] make use of the supervised contrastive loss function. Although the contrastive learning strategy is deemed feasible to achieve the class alignment stage, it requires the sample diversity, which is not served by the naïve data augmentation strategy, such as the Gaussian noise. Besides, we offer a new perspective in this paper where the PEFT strategy is proposed for the hyperspectral foundation model.
III Preliminaries
III-A Problem Definition
@@
III-B Mixup Strategy
Our approach, MIFOMO, leverages the mixup strategy to enhance the sample’s diversity, deemed as the consistency-based regularization strategy, forcing the network to output the same predictions for interpolated samples [31]. It has been shown to meet a lower bound of the Lipschitz constant of the gradient of the neural network. In the realm of the HSI classification problem using the deep learning approach, this strategy can be committed in both the image level or the embedding level and generates an augmented sample pair based on a linear interpolation strategy. Given two distinct HSIs $(x_{i},x_{j})$ and their corresponding labels $(y_{i},y_{j})$, the mixup method is described as follows:

### v2

@@ Abstract
Remote Sensing, Cross-Domain Few-Shot Learning, Few-Shot Learning, Meta-learning, Transfer Learning, Hyperspectral Image, Sustainable Development Goal (SDG).
1 Introduction
@@
A Hyperspectral image (HSI) plays an important role in the remote sensing (RS) field because it is, unlike RGB images, in the form of 3D data cubes and contains abundant spatial and spectral information revealing unique characteristics of various land covers Landgrebe2002HyperspectralID. It is widely applied across different application domains such as mining, biology, and natural disasters, etc.
@@
The challenge of the HSI classification task is to assign a label to each pixel, which incurs high labeling costs. Early approaches to HSI classification problems make use of traditional machine learning methods Samaniego2008SupervisedCO; Melgani2004ClassificationOH combined with manual feature extraction techniques. Nonetheless, such approaches are over-dependent on prior knowledge and suffer from significant bottlenecks in handling high-dimensional data such as HSIs. The advent of deep learning (DL) has revolutionized the RS field and has been successful across different RS problems, including the HSI classification problems. In contrast to traditional approaches, the DL methods feature an automatic feature engineering process, making it possible for deep and nonlinear features to be automatically extracted, learned, and classified Chen2014DeepLC. Various models have been proposed to cope with the HSI classification problems, including CNN Lee2016GoingDW; Zhao2022SuperpixelGD; Zhang2018DiverseRC, transformer Hong2021SpectralFormerRH; Wang2023DCNTDC, graph neural network Liu2020CNNEnhancedGC; Dong2022WeightedFF, etc. Nonetheless, the success of DL approaches is attributed to the availability of large datasets hindering their practical applications in the real-world HSI classification problems, imposing prohibitive labeling costs Paeedeh2024FewShotCI.
@@
Such an issue has motivated the development of various methods to reduce the labeling cost problem. The semi-supervised learning (SSL) Yao2023PseudolabelBasedUS; Wei2023SemiSupervisedNA and self-supervised learning (SSL) Li2024CrossDomainFH; Ma2023SelfSupervisedSG methods have been proposed in the realm of the HSI classification problems. Although these approaches are capable of reducing the labeling cost to some extent, they call for a lot of samples. It is not well-suited to the HSI classification problem, where the cost of data collection is also expensive. This issue has led to the development of few-shot learning methods Snell2017PrototypicalNF; Sung2017LearningTC for the HSI classification problems. The few-shot learning approach relies on an episodic learning procedure based on the so-called meta-learning approach. That is, the training process is divided into a number of episodes having very few samples, i.e., support and query sets. Such approaches allow to quickly generalize to unseen classes with limited samples because a model is trained to acquire meta-knowledge. Nevertheless, conventional few-shot learning methods aren’t prepared to address the domain shift problem where the source and target domains are drawn from different distributions. This bottleneck is not in line with the fact that data collection of HSIs might be influenced by external factors, e.g., weather, lighting, sensor nonlinearities, etc., causing the issue of domain shifts.
@@
Cross-scene HSI classification methods Yang2016DomainAW; Ma2021CrossDatasetHI; Huang2022TwoBranchAA are proposed to address the domain shift problem. This approach transfers knowledge of a label-rich domain known as the source domain to a label-scarce domain known as the target domain. The underlying assumption is that there are no labels at all in the target domain, and the presence of domain shifts between the source domain and the target domain. Usually, such approaches are driven by the domain adaptation (DA) method Huang2023CrosssceneWM to align the distributional discrepancies between the source domain and the target domain. The key limitation of the cross-scene HSI classification lies in the closed-set assumption, where the source domain and the target domain share the same label space. Cross-Domain Few-Shot Learning (CDFSL) Zhang2022GraphIA; Paeedeh2024CrossDomainFL aims to overcome this limitation, where the goal is to cope with the issue of data scarcity and domain shift simultaneously. It goes one step ahead of cross-scene HSI classification method,s where the label space between the source domain and the target domain can be completely different.
@@
The CDFSL Paeedeh2024CrossDomainFL for HSI classification has grown rapidly, where numerous works have been proposed to deal with this problem. Existing works can be categorized into three groups: discrepancy-based methods Li2024SCFormerSC; Hu2023CrossDomainMU; Li2024SemanticGP; Dang2025CrossDomainFL; Qin2024FewShotLW; Zhang2022GraphIA; Qin2024CrossDomainFL; Feng2024CrossDomainFL, adversarial-based methods Peng2023ConvolutionalTF; Shi2025MultibranchFT; Qin2024FewShotLW; Wang2022SpatialSpectralLD; Zhang2023CrossDomainSN; Ye2023AdaptiveDF; Cao2024SpatialSpectralSemanticCF; Wang2024DualBranchDA; Liu2025MultilevelPA; Shi2025MultiGaussianPM; Zhu2025FromIT and contrastive-based methods Shi2025FewShotLB; Ye2024CrossDomainFL; Huang2023HyperspectralIC; Zhang2022CrossDomainFC; Li2024CrossDomainFH. The discrepancy-based methods are designed to minimize distributional discrepancy between the source domain and the target domain using the discrepancy metrics. The adversarial-based methods utilize the adversarial domain adaptation technique Ganin2015DomainAdversarialTO, playing the minmax game between the discriminator and generator. The contrastive-based methods apply the idea of contrastive learning to guarantee the compactness of intra-class data and the separability of inter-class data. Nonetheless, existing methods suffer from at least three drawbacks:
@@
  • Existing methods apply unrealistic data augmentation procedures to enlarge the size of the target domain Wang2024DualBranchDA. Such an approach greatly simplifies the issue of data scarcity but may be trapped in the out-of-distribution problem. The model performance may be compromised by the problem of sample diversity. In addition, the data augmentation procedure simply follows the RGB image augmentation strategy. The RGB image augmentation strategy might not be compatible with the HSI case.
@@
  • Very few works have explored the strength of hyperspectral foundation models in the realm of CDFSL for HSI classification problems. The foundation model possesses strong generalization potential, which can be quickly adapted to downstream tasks, thus aligning very well with the spirit of the CDFSL for HSI classification problems. To the best of our knowledge, only Chen2025SpectralDINODM has pioneered the use of a foundation model in the context of CDFSL for HSI classification. However, this method still utilizes the foundation model of the RGB images, hindering their generalization for HSIs. Besides, we propose a new parameter-efficient fine-tuning (PEFT) method, namely coalescent projection (CP), whereas they Chen2025SpectralDINODM are still based on the LoRA method.
This paper proposes a mixup foundation model (MIFOMO) for CDFSL of HSI classification problems. MIFOMO is built upon a hyper-spectral foundation model, HyperSIGMA Wang2024HyperSIGMAHI, pretrained in the HyperGlobal-450K dataset containing 450 K HSIs. This goes one step ahead of Chen2025SpectralDINODM, still relying on the foundation model of RGB images, dinoV2. To remedy the issue of overfitting, we put forward a novel PEFT technique, namely CP Paeedeh2025CrossDomainFL. That is, only CPs are trainable during the training process, leaving the backbone network frozen. This strategy also allows preservation of the strong generalization power of the foundation model. CP can be seen as a successor of plain prompt Wang2021LearningTP, incurring significantly fewer parameters because it relies on a single learnable matrix. In addition, the CP technique is simpler to apply than the plain prompt because there is no need to select the prompt length. Another innovation of MIFOMO lies in the domain adaptation strategy, where the mixup domain adaptation approach Furqon2024MixupDA is put into perspective. This strategy consists of two phases: source domain and intermediate domain, in which the episodic-based meta-learning approach is undertaken. Unlike direct knowledge transfer between the source and target domains, the intermediate domain functions as a bridge between the source and target domains, making seamless knowledge transfer possible because direct domain alignment may be too intricate to perform. Last but not least, the concept of label smoothing is incorporated to combat the noisy pseudo-label problem. That is, pseudo-labels are assigned to the query set of the target domain to prevent the label leakage problem.
@@
    This paper proposes the mixup foundation model (MIFOMO) for the CDFSL of HSI classification problems. This approach is unique in respect to state-of-the-art (SOTA) methods because it relies on the PEFT approach, i.e., existing works are based on the discrepancy-based methods, the adversarial-based methods, and the contrastive-based methods. It distinguishes itself from Chen2025SpectralDINODM because of the hyperspectral foundation model and the CP technique.
@@
    The concept of CP is introduced to adapt the foundation model, namely HyperSIGMA, to the target domain. CP uses a single trainable matrix connecting the key and query projections in the attention map while leaving the backbone network frozen. CP offers very few trainable parameters and is easy to apply because it isn’t dependent on any hyperparameters. We extend Paeedeh2025CrossDomainFL designed for natural images to hyperspectral images here.
@@
    MIFOMO advances the mixup domain adaptation Furqon2024MixupDA originally designed for unsupervised domain adaptation (UDA) scenarios, i.e., our concern is in the CDFSL of the HSI classification problem here. Moreover, the mixup domain adaptation method is generalized here using the idea of label smoothing to combat the noisy pseudo-label problem when meta-learning the target domain, while the episodic-based meta-learning approach is implemented in the source and intermediate domains. Besides, Furqon2024MixupDA still works on the traditional architecture, whereas we implement such an idea in the context of a foundation model.
@@ 2 Related Works
Cross-Domain Few-Shot Learning (CDFSL) for the Hyperspectral Image (HSI) classification problem is meant to address the problem of data scarcity and domain shift in the HSI classification problem. It aims to leverage discriminative information of the label-rich domain (source domain) to solve the HSI classification problem within the label-scarce domain (target domain). Compared to cross-scene HSI classification problems Yang2016DomainAW; Ma2021CrossDatasetHI; Huang2022TwoBranchAA,this problem features extreme domain discrepancies where the label spaces between the two domains might be completely disjoint. The CDFSL for HSI classification problems have grown rapidly where existing approaches can be divided into three categories based on how the domain alignment approach is performed: discrepancy-based methods Li2024SCFormerSC; Hu2023CrossDomainMU; Li2024SemanticGP; Dang2025CrossDomainFL; Qin2024FewShotLW; Zhang2022GraphIA; Qin2024CrossDomainFL; Feng2024CrossDomainFL, adversarial-based methods Peng2023ConvolutionalTF; Shi2025MultibranchFT; Qin2024FewShotLW; Wang2022SpatialSpectralLD; Zhang2023CrossDomainSN; Ye2023AdaptiveDF; Cao2024SpatialSpectralSemanticCF; Wang2024DualBranchDA; Liu2025MultilevelPA; Shi2025MultiGaussianPM; Zhu2025FromIT and contrastive-based methods Shi2025FewShotLB; Ye2024CrossDomainFL; Huang2023HyperspectralIC; Zhang2022CrossDomainFC; Li2024CrossDomainFH.
2.1 Discrepancy-based method
@@
The discrepancy-based method measures the domain gap and minimizes it via a distance metric. Li2024SCFormerSC utilizes the Wasserstein distance metric, Hu2023CrossDomainMU applies the class covariance matrix, Li2024SemanticGP uses the semantic information as complementary information of prototype construction, Dang2025CrossDomainFL implements the CORAL strategy, Qin2024FewShotLW makes use of the MMD metric. Zhang2022GraphIA applies the graph mining strategy, Qin2024CrossDomainFL designs the multiorder spectral interaction block, Feng2024CrossDomainFL uses the knowledge distillation strategy. The aforementioned works directly minimize the discrepancy between two domains, which can be difficult in the case of extreme domain gaps and often result in negative transfer. Our approach here is based on the mixup domain adaptation, consisting of two stages: the source domain and the intermediate domain. The intermediate domain is crafted through the mixup mechanism and allows smooth knowledge transfer because it functions as a bridge between the source domain and the target domain. In addition, the direct domain adaptation method performs global domain alignments, ignoring the class alignment step, i.e., local. Note that the label spaces of the two domains don’t overlap.
2.2 Adversarial-based Method
@@
The adversarial-based method follows an adversarial domain adaptation strategy Ganin2015DomainAdversarialTO which introduces a domain discriminator detecting the origin of data samples, i.e., source or target domains. It plays a minmax game with the feature extractor, and the domain alignment stage is achieved when the feature extractor successfully fools the domain discriminator. Peng2023ConvolutionalTF applies the vanilla adversarial domain adaptation method, Shi2025MultibranchFT implements the conditional adversarial domain adaptation method, Qin2024FewShotLW constructs an intermediate domain, Wang2022SpatialSpectralLD devises the local and global adversarial domain adaptation method, Zhang2023CrossDomainSN utilizes the collaborative adversarial network concept. Ye2023AdaptiveDF inserts a varying weight to the conditional adversarial loss, Cao2024SpatialSpectralSemanticCF uses a semantic-aware strategy for the domain discriminator, Wang2024DualBranchDA implements the random sampling strategy to address the curse of dimensionality, and Liu2025MultilevelPA applies the prototype-based conditional adversarial domain adaptation method. Shi2025MultiGaussianPM proposes a cyclic domain triplet loss combining the adversarial domain adaptation technique and the contrastive learning strategy, while Zhu2025FromIT offers a cyclic resemblance domain adversarial network. As with the discrepancy-based method, the adversarial-based method performs the global domain alignment strategy, excluding the class alignment step.
2.3 Contrastive-based method
@@
The contrastive-based method relies on the contrastive learning strategy to attain alignments between the two domains. Shi2025FewShotLB proposes the inter-domain contrastive loss function, Ye2024CrossDomainFL puts forward the graph convolutional contrast module, Huang2023HyperspectralIC puts into perspective the kernel triplet loss function, Zhang2022CrossDomainFC; Li2024CrossDomainFH make use of the supervised contrastive loss function. Although the contrastive learning strategy is deemed feasible to achieve the class alignment stage, it requires the sample diversity, which is not served by the naïve data augmentation strategy, such as the Gaussian noise. Besides, we offer a new perspective in this paper where the PEFT strategy is proposed for the hyperspectral foundation model.
3 Preliminaries
3.1 Problem Definition
@@
3.2 Mixup Strategy
Our approach, MIFOMO, leverages the mixup strategy to enhance the sample’s diversity, deemed as the consistency-based regularization strategy, forcing the network to output the same predictions for interpolated samples Mai2019MetaMixUpLA. It has been shown to meet a lower bound of the Lipschitz constant of the gradient of the neural network. In the realm of the HSI classification problem using the deep learning approach, this strategy can be committed in both the image level or the embedding level and generates an augmented sample pair based on a linear interpolation strategy. Given two distinct HSIs $(x_{i},x_{j})$ and their corresponding labels $(y_{i},y_{j})$, the mixup method is described as follows: